0 0
Add a comment Improve this question Transcribed image text
Answer #1

Part B:

The light is shone on silver, photo electrons emerge eith zero velocity. So the energy of photon which is absorbed by an electron of silver, is just enough to release it from the surface of silver. In others words the energy of a incident photon is equal to the work function of silver = 4.73 eV.

Same light when shone on sodium, whose work function is 2.46 eV. As light is always absorbed in integer multiples, the photon cannot eject two photo electrons from surface. So when an electron absorbs a photon, the excess energy of photon is observed as kinetic energy of the electron= 4.73eV - 2.46eV= 2.27eV

Part C:

The energy of a nth stationary level in a hydrogen atom, is given by

E_n=[\frac{2\pi^2 me^4}{h^2}]\ \frac{1}{n^2}=\frac{-13.6}{n^2}\ eV

where m and e are mass and charge of an electron and h is planck's constant

The magnitude of energy of a stationary level decreases, with increasing n. Pictorially

A dexcitation is a phenomena in which an electron moves from high energy level (excited state) to low energy state. So we the number of transitions would be the equal to ways to select any two levels out of 3 total levels.

Thus we have 3_{C_2} = 3 transitions totally, from n=3 to n=2 shown as red ; n=3 to n=1 shown as blue ; n=2 to n=1 shown as green

E_{photon\ in\ eV}X q_{electron}=E_{photon\ in\ joule}=\frac{hc}{\lambda}\Rightarrow \lambda=\frac{hc}{qE_{ev}}

This a handy conversion formula \lambda_{A^o}=\frac{12406}{E_{ev}}\ and\ 1A^o=10nm

The red code line emission has an energy =1.89 eV which corresponds to 656.4nm

The green coded line emission has an energy= 10.20 eV which corresponds to 121.6nm

The blue coded line emission corresponds to= 12.09 eV which corresponds to 102.6nm
Answer #2

b) When the electrons are emitted with almost zero velcoity the work function is equal to the incident energy. so incident energy is 4.73 eV

now using einstein photoelectric equation

        E = W + K

so kinetic energy k = E - W = 4.73 - 2,46 = 2.27 eV

C) II)

The three possible transistions are n=3 to n=2, n=3 to n=1 & n=2 to n=1.

the wavelength equation is 1/\lambda = R [ (1/n32) - (1/n22) ] = 656.3nm

                                       1/\lambda = R [ (1/n32) - (1/n12) ] = 102.6nm

                                       1/\lambda = R [ (1/n22) - (1/n12) ] = 121.6nm